# Evaluating and improving wave and non-wave stress parametrisations for oceanic flows

### v1

@@ Abstract
Whenever oceanic currents flow over rough topography, there is an associated stress that acts to modify the flow. In the deep ocean, this stress is predominantly form drag, caused by the formation of internal waves and other baroclinic motions: processes that act on such small scales most global ocean models cannot resolve. Despite the need to incorporate this stress into ocean models, existing parametrisations are limited in their applicability. For instance, most parametrisations are only suitable for small-scale topography and are either for periodic or steady flows, but rarely a combination thereof. Here we summarise these existing parametrisations and evaluate their accuracy using hundreds of idealised two-dimensional and three-dimensional simulations spanning a wide parameter space. In cases where the parametrisations prove to be in inaccurate, we use our data to suggest improved formulations. Our results thus provide a basis for a comprehensive parameterisation of stress in ocean models where fine scale topography is unresolved.
@@
As the ocean flows over topographic features it experiences a force. Such topographic features include abyssal hills that are too small to be accurately described by global ocean and climate models. These forces shape the ocean structure considerably and affect large-scale ocean circulation features, like the Atlantic Meridional Overturning Circulation. Because there is no hope to accurately include these small-scale features in climate projections we need expressions (a.k.a. parametrisations) for the collective force they exert on the ocean. Here, we evaluate existing parametrisations for the force created when an oceanic current flows over an underwater hill by comparing them to the output of small-scale, high-resolution ocean simulations. Compared to previous studies, we cover a much wider range of cases. This allows us to suggest improvements to the existing parametrisations, taking us closer to the goal of developing a general and widely applicable mathematical description of this phenomenon.
@@ 1 Introduction
[…]
@@
Due to the complexity of developing a generic parametrisation for stress, existing parametrisations (e.g., Bell, 1975; Jayne and St. Laurent, 2001; Klymak et al., 2010; Shakespeare et al., 2020) have focused on special cases. For instance, previous studies have treated steady and tidal flows in isolation. However, since the steady and tidal components of a flow can interact (Shakespeare, 2020), it is important for existing parametrisations to be extended to mixed flows. Existing parametrisations also have several other assumptions that can reduce their accuracy if incorporated into ocean models. For example, it is often assumed that the topographical height is small (Bell, 1975; Shakespeare et al., 2020) or that the latitude is close to $0^{\circ}$ (Jayne and St. Laurent, 2001; Klymak et al., 2010).
@@
The first goal of this paper is to evaluate the accuracy of existing wave and non-wave stress parametrisations. This is done by comparing theoretical expressions to a large suite of ocean simulations. To limit the scope of these comparisons, we focus on the case of an isolated Gaussian hill in either two or three dimensions. Working with isolated hills is useful as there are only a small number of parameters to consider, allowing us to ensure that the basic physics and scalings of the parametrisations are correct. This is in contrast to other studies (e.g. Jayne and St. Laurent (2001); Buijsman et al. (2015); Shakespeare et al. (2021)) which tend to only test stress parametrisations with global models or more complicated topography. We also note that whilst three-dimensional simulations are physically accurate, they are more computationally expensive than two-dimensional simulations. Moreover, two-dimensional topography may be interpreted as approximating a section of a three-dimensional ridge, which can be sites of significant internal wave generation (Ledwell et al., 2000; Garrett and Kunze, 2007). Two-dimensional simulations are thus useful despite being less applicable than three-dimensional models. As a result, our general procedure is to run a large number of tests in two dimensions, and then further analyse any interesting phenomena in three dimensions.
@@
The second goal of this paper is to then suggest improvements to these existing stress parametrisations. This is done using predominantly mathematical or physical arguments that are supported by the simulation data. Notably, one of the common shortcomings in existing parametrisations is the assumption of small topographical height. Large topographical features however, are abundant in the ocean (Harris et al., 2014). Moreover, large isolated hills called seamounts are of interest in ecology, vulcanology and ship navigation (Clark et al., 2010; Watts, 2019). Therefore, it is important for existing parametrisations to be extended to such cases. The ultimate goal of this work is the formulation of a more complete parametrisation of wave and non-wave stress for implementation in global ocean models.
@@
The structure of the paper is as follows. In Section 2 we recall some existing drag parametrisations used for tidal, steady and mixed flows. In Section 3 we discuss the setup for our simulations. In Section 4 we then discuss the results of the simulations, and use these results to evaluate and suggest improvements to each parametrisation. Finally, in Section 5 we discuss the main takeaways from our analysis, and prescribe new parametrisations that depend on the type of flow and scale of topography under consideration.
@@ 2 Existing parametrisations
We begin by summarising existing parametrisations for wave and non-wave stress. The parametrisations are categorised depending on the flow in consideration — tidal, steady or mixed. Each parametrisation is expressed in their general form, but we also simplify for the case of an isolated Gaussian hill. For a two-dimensional domain (viewed as cross-section of three dimensional space) with horizontal component $[-L_{x}/2,L_{x}/2]$, we define the stress as
@@
where $h(x)$ is the bottom topography, $p_{h}(x)$ the bottom pressure respectively, and $\rho_{0}$ is the background density of the sea water. In particular, the stress (1) is the net force per unit length [^1] That is, per unit length in the $y$-direction. Hence, the two-dimensional stress can be used to compute the total force across a long three-dimensional ridge. created from the pressure differential across the hill, divided by $\rho_{0}$ for convenience. As a result, $F_{\text{2d}}$ has units $\mathrm{N}\,\mathrm{m}^{2}\,\mathrm{k}\mathrm{g}^{-1}$ and $\rho_{0}F_{\text{2d}}$ has units of force per unit length. The sign of $F_{2d}$ is also such that a positive value of $F_{2d}$ corresponds to a force per unit length of magnitude $|\rho_{0}F_{\text{2d}}|$ acting on the hill in the positive $x$-direction, along with an equal and opposite force acting on the flow in the negative $x$-direction. In three dimensions, with a horizontal domain $[-L_{x}/2,L_{x}/2]\times[-L_{y}/2,L_{y}/2]$, we then define the stress due to a topography $h(x,y)$ as the vector
@@
Note that $\bm{F}_{\text{3d}}$ has units $\mathrm{N}\,\mathrm{m}^{3}\,\mathrm{k}\mathrm{g}^{-1}$ and $\rho_{0}\bm{F}_{3d}$ has units of force (i.e., $\mathrm{N}$).
In what follows, our domain is as shown in Figure 1, with dimensions $L_{x}\times L_{y}\times H$ and a spatially mean zonal flow $U(t)$ over topography $h(x,y)$. For two-dimensional analysis, we restrict ourselves to the $x-z$ plane. Moreover, for the ensuing theoretical analysis in this section, we assume $L_{x}$ and $L_{y}$ are large. That is, $L_{x},L_{y}\to\infty$.
@@
Throughout this paper, we use the terms “stress”, “form drag” and “internal wave drag” somewhat interchangeably. However, it is to be understood that each of these terms are defined differently. By “stress” we mean the total force (scaled by $1/\rho_{0}$) acting on the current as per (1) or (2). Then, by “form drag” we mean the component of the stress in phase with the current and therefore associated with a mean energy exchange. Finally, “internal wave drag” refers to the component of form drag due to the formation of internal waves. Many existing parametrisations focus only on internal wave drag, which is the dominant component of stress in most idealised scenarios, but not if the flow is “bottom-trapped”, sufficiently non-linear or resonant with the topography.
@@ 2.1 Case 1: Tidal Flow
We consider a periodic tidal flow $U(t)=U_{\text{tidal}}\cos(\omega t)$ in the $x$-direction where $U_{\text{tidal}}$ is a constant and $\omega$ is the tidal frequency. Tidal flows are a major source of waves and the associated stress impacts the strength of barotropic and internal tides (Ansong et al., 2015; Buijsman et al., 2015). As a result, this flow regime appears to be the main focus of stress in the literature. Here, we restrict our focus to two parametrisations. Namely, an older parametrisation due to Jayne and St. Laurent (2001, hereafter, JSL2001), and a more recent parametrisation due to Shakespeare et al. (2020, hereafter SAH2020).
@@ 2.1.1 Jayne and St Laurent parametrisation
Jayne and St Laurent give a simple estimate for internal wave drag based on a scaling argument derived from the work of Bell (1975). Namely, for each grid cell, they set the stress (in each horizontal dimension) to be proportional to $Nh^{2}u$. Here, $N$ is the buoyancy frequency, and $h$ and $u$ are taken to be suitable values of the topography height and tidal velocity in the grid cell. For small $h\ll H$, we can approximate $u$ as $U(t)$ and $N$ as a constant in the vicinity of the topography. Thus, the total stress over a small hill with root mean square height $h_{rms}$ is estimated by
@@
Here, $\kappa$ is the tunable scaling constant in JSL2001 related to the characteristic wavenumber of the topography. Then, $N_{B}$ denotes the buoyancy frequency at the seabed, and the subscript “JSL” indicates that this is the drag force due to the JSL2001 theory. In line with our units and scaling for stress in (1) and (2), here $\kappa$ has units of length in three dimensions and is unitless in two dimensions. On the other hand, in JSL2001 the stress is implicitly divided by the horizontal length (in two-dimensions) or area (in three-dimensions) so that $\kappa$ has units of inverse length. Note that the factor of $1/2$ in (3) could be absorbed into the definition of $\kappa$, but we have included it to more closely match the expression given by JSL2001.
Despite the simple form of the JSL2001 parametrisation, this scaling has proven to significantly improve the accuracy of tidal models when compared to observations, and is thus frequently used today (Arbic et al., 2018; Ansong et al., 2015). However, there are some notable limitations of the JSL2001 model. Firstly, the model is only designed for small-scale topography and large ocean depths. Moreover, (3) is latitude-independent, an omission which Jayne and St Laurent state could be significant when the tidal frequency $\omega$ is close to the Coriolis frequency $f$.
@@ 2.1.2 Shakespeare, Arbic and Hogg parametrisation
In SAH2020, the authors take a more analytic approach to parametrising tidal flow stress. The main novelty of their work is their treatment of “spring” forces, such as bottom-trapped tides, which do not do any work in the time mean. This is important, because JSL2001 and other early parametrisations compute stress by first considering the time-averaged energy dissipation, and thus do not account for such spring forces. Under standard assumptions, including small topographical height and large ocean depth, SAH2020 give for subcritical latitudes ($|f|<\omega$) and isotropic topography
@@
Here, $\hat{h}$ denotes the Fourier transform of $h$ in $t$, $x$ and $y$ to transformed variables $\omega$, $k$ and $\ell$ respectively with $K=\sqrt{k^{2}+\ell^{2}}$. Note that we use the Fourier transform convention
@@
Since we are assuming the topography is isotropic (and independent of time), $\hat{h}$ is a function of the total wavenumber $K$ only, as indicated in (4). We have also introduced a parameter $\alpha\in\{0,1\}$, whereby
In addition, we remark that (4) is only the $x$-component of the stress, as this is the primary direction of our flow and focus of our analysis.
In SAH2020, a more general expression than (4) is stated which accounts for wave reflections off the surface. However, such reflections are difficult to parametrise, and for an isolated hill the number of waves reflecting back onto the topography is small. Therefore, we omit any reflection terms from (4) for simplicity. Nevertheless, we note that these reflections depend on how the vertical wavelength
$m=\frac{Nk}{\sqrt{\omega^{2}-f^{2}}},$ (7)
resonates with the ocean depth $H$. As a result, we expect (4) to have small deviations from the actual stress as the geometry of the topography, and the variables $N$, $\omega$ and $f$ are varied.
In the case of an isolated Gaussian with half-width $W$, whereby
@@
Note that in the limit $N\gg\omega$ (or $\alpha=0$) and $\omega\gg|f|$, (9) is of the same form as (3). The expression for supercritical (bottom-trapped) latitudes $|f|\geq\omega$ is similar, but with the stress out of phase with velocity. In particular, for a Gaussian hill
@@
×$\ \times\begin{cases}\sqrt{(N^{2}-\alpha\omega^{2})(\omega^{2}-f^{2})}\,\cos(\omega t),\ |f|<\omega,\\ \sqrt{(N^{2}-\alpha\omega^{2})(f^{2}-\omega^{2})}\,\sin(\omega t),\ |f|\geq\omega,\end{cases}$ (10)
where the subscript SAH3d indicates that this is the three-dimensional stress due to the SAH2020 theory. In a subsequent paper (Shakespeare et al., 2021), the authors also give a parametrisation for two dimensions. Namely, for a Gaussian hill $h(x)=h_{0}e^{-x^{2}/(2W^{2})}$,
@@
×$\ \times\begin{cases}\sqrt{(N^{2}-\alpha\omega^{2})(\omega^{2}-f^{2})}\cos(\omega t),&|f|<\omega,\\ \sqrt{(N^{2}-\alpha\omega^{2})(f^{2}-\omega^{2})}\sin(\omega t),&|f|\geq\omega.\end{cases}$ (11)
Note that (2.1.2) is identical to (2.1.2) except without the factor of $\pi\sqrt{\pi}W/4$. To obtain (2.1.2) and (2.1.2) one also needs to assume that $N$ is constant. In the more physical setting where $N$ is varying function of depth, Shakespeare et al. (2021) suggest taking $N$ to be the value of the buoyancy frequency at the mean ocean depth.

### v3

@@ Abstract
Whenever oceanic currents flow over rough topography, there is an associated stress that acts to modify the flow. In the deep ocean, this stress is predominantly a form drag due to pressure differentials across topography, caused by the formation of internal waves and other baroclinic motions: processes that act on such small scales most global ocean models cannot resolve. Despite the need to incorporate this stress into ocean models, existing parametrisations are limited in their applicability. For instance, most parametrisations are only suitable for small-scale topography and are either for periodic or steady flows, but rarely a combination thereof. Here we summarise some of the most widely used parametrisations and evaluate the accuracy of a carefully selected subset using hundreds of idealised two-dimensional and three-dimensional simulations spanning a wide parameter space. We focus on the case of an isolated Gaussian hill as an idealised representation of a seamount. In cases where the parametrisations prove to be inaccurate, we use our data to suggest improved formulations. Our results thus provide a starting point for a comprehensive parameterisation of topographic stresses in ocean models where fine scale topography is unresolved.
@@
As the ocean flows over topographic features it experiences a force. Such topography includes undersea mountains that are too small to be accurately described by global ocean and climate models. These forces shape the ocean structure considerably and affect large-scale ocean circulation features, like the Atlantic Meridional Overturning Circulation. Because there is no hope to accurately include these small-scale effects in climate projections we need expressions (a.k.a. parametrisations) for the collective force they exert on the ocean. Here, we evaluate existing parametrisations for the force created when an oceanic current flows over an underwater hill by comparing them to the output of small-scale, high-resolution ocean simulations. Compared to previous studies, we cover a much wider range of cases. This allows us to see where existing parametrisations break down and also suggest improvements to these parametrisations, taking us closer to the goal of developing a general and widely applicable mathematical description of this phenomenon.
@@ 1 Introduction
Over half of the ocean floor is covered by abyssal hills and mountains, defined respectively as having a variation in height of 3001000300-1000 m and greater than 1000m1000\ $\text${m} (Harris et al., 2014) over a 750 arc second (or $\approx 10-20\ \mathrm{km}$) radius. Since internal waves generated at these features are of a comparable horizontal scale to the topography itself (e.g., Llewellyn Smith and Young, 2002), resolving these waves is not possible with current global ocean models which generally have horizontal resolutions on the order of 10 to 100 km (e.g., Ansong et al., 2015; Kiss et al., 2020). As a result, it is important to have parametrisations for the effects of internal waves generated at such hills. Existing parametrisations have mainly focused on computing the energy flux extracted from the flow due to the generation of internal waves (e.g., Simmons et al., 2004; Jayne, 2009; Olbers and Eden, 2013; de Lavergne et al., 2019). Such a focus is justified, as it has allowed better understanding of the contribution of internal waves to tidal and eddy dissipation, and ocean mixing. However, in recent times, there have been further efforts (e.g., Shakespeare et al., 2020; Klymak et al., 2021; Eden et al., 2021) which attempt to develop parametrisations for the corresponding topographic stress due to internal waves and other related effects. These effects include blocking (Winters and Armi, 2014; Nikurashin et al., 2014) and bottom-trapped tides formed at high latitudes where a stronger Coriolis force prevents internal waves from freely propagating (Falahat and Nycander, 2015; Shakespeare et al., 2020). Developing these stress parametrisations is vital to obtaining accurate representations of eddies and tides, and large-scale flow features such as the Antarctic Circumpolar Current (e.g., Naveira Garabato et al., 2013; Arbic et al., 2019).
To be precise, the energy loss from a spatially uniform flow $\overline{\bm{u}}$ and topographic stress $\bm{\tau}$ is equal to $\overline{\bm{u}}\cdot\bm{\tau}$, and the topographic stress also acts to modify $\overline{\bm{u}}$. Thus, the topographical stress is a more fundamental quantity to study, in that it can be used to compute energy loss, but this is not necessarily true the other way around. This is particularly the case when $\overline{\bm{u}}$ is time-varying, so that it is not at all valid to use expressions for the time-averaged energy flux to understand the stress $\bm{\tau}$. Doing so omits any component of the stress $\bm{\tau}$ that is out-of-phase with the flow $\overline{\bm{u}}$, which in some cases is the entire stress (e.g., for bottom-trapped tides; Falahat and Nycander, 2015; Shakespeare et al., 2020)
@@
Due to the complexity of developing a generic parametrisation for stress, existing work (e.g., Bell, 1975; Jayne and St. Laurent, 2001; Klymak et al., 2010; Shakespeare et al., 2020) has focused on special cases. For instance, previous studies have treated steady and tidal flows in isolation. However, since the steady and tidal components of a flow can interact (Shakespeare, 2020), it is important for existing parametrisations to be extended to mixed flows. Existing parametrisations also have several other assumptions that can reduce their accuracy if incorporated into ocean models. For example, it is often assumed that the topography height is small (Bell, 1975; Shakespeare et al., 2020) or that the latitude is close to $0^{\circ}$ (Jayne and St. Laurent, 2001; Klymak et al., 2010).
@@
The first goal of this paper is to evaluate the accuracy of a subset of available wave and non-wave stress parametrisations. This is done by comparing theoretical expressions to the stress calculated from a large suite of ocean simulations. The main parametrisations we focus on are those due to Bell (1975), Jayne and St. Laurent (2001), Klymak et al. (2010) and Shakespeare et al. (2020), and we refer to the reader to other related works throughout. Here we focus on the magnitude of the stress the fluid exerts on the topography, rather than discerning exactly where in the water column the equal and opposite reaction stress on the ocean is applied. We also focus on the case of an isolated Gaussian hill in either two or three dimensions, which is a standard, simple case which has been studied in the context of related phenomena (Legg and Huijts, 2006; Maitland-Davies and Bühler, 2025). A Gaussian hill is a decent model for an isolated undersea mountain (seamount). However, a stronger motivation in this case is that working with isolated hills means that there are only a small number of well-defined parameters to consider (i.e., height and width), allowing us to ensure that the basic physics and scalings of the parametrisations are correct. This is in contrast to other studies (e.g., Jayne and St. Laurent, 2001; Buijsman et al., 2015; Shakespeare et al., 2021a) which tend to only test stress parametrisations with global models or more complicated topography, and often with more limited parameter regimes. In a similar vein, we reduce the complexity of our simulations by taking the stratification to be constant. We also note that whilst three-dimensional simulations are more physically realistic, they are more computationally expensive than two-dimensional simulations. Moreover, two-dimensional topography may be interpreted as approximating a section of a three-dimensional ridge, which can be sites of significant internal wave generation (Ledwell et al., 2000; Garrett and Kunze, 2007). Two-dimensional simulations are thus useful despite being less broadly applicable than three-dimensional models. As a result, our general procedure is to run a large number of tests in two dimensions, and then further analyse any interesting phenomena in three dimensions.
@@
The second goal of this paper is to suggest improvements to existing stress parametrisations. This is done using predominantly mathematical or physical arguments that are supported by the simulation data. Notably, one of the common shortcomings in existing parametrisations is the assumption of small topographic height despite tall topographic features being abundant in the ocean (Harris et al., 2014). Moreover, large isolated seamounts are of interest in ecology, vulcanology and ship navigation (Clark et al., 2010; Watts, 2019). Therefore, it is important for existing parametrisations to be extended to such cases. The ultimate goal of this work is to understand the isolated hill regime in detail and help pave the way for a more complete parametrisation of wave and non-wave stress for implementation in global ocean models.
@@
The structure of the paper is as follows. In Section 2 we recall some existing drag parametrisations used for tidal, steady and mixed flows. In Section 3 we discuss the setup for our simulations. In Section 4 we then discuss the results of the simulations, and use these results to evaluate and suggest improvements to each parametrisation. Finally, in Section 5 we discuss the main takeaways from our analysis, and describe a combined parametrisation that depends on the type of flow and scale of topography under consideration.
@@ 2 Existing parametrisations
We begin by summarising existing parametrisations for wave and non-wave stress. The parametrisations are categorised depending on the flow under consideration — tidal, steady or mixed. Each parametrisation is expressed in its general form, but we also simplify for the case of an isolated Gaussian hill. For a two-dimensional domain (viewed as cross-section of three-dimensional space) with horizontal component $L_{x}/2,L_{x}/2]$, we define the stress as
@@
where $h(x)$ is the bottom topography, $p_{h}(x)$ the bottom pressure, and $\rho_{0}$ is the background density of the sea water. In particular, the stress (1) is the net force per spanwise unit length created from the pressure differential across the hill, divided by $\rho_{0}$ for convenience. As a result, $F_{\text{2d}}$ has units $\mathrm{N}\,\mathrm{m}^{2}\,\mathrm{k}\mathrm{g}^{-1}$ and $\rho_{0}F_{\text{2d}}$ has units of force per unit length. The sign of $F_{\text{2d}}$ is also such that a positive value of $F_{\text{2d}}$ corresponds to a force per unit length of magnitude $|\rho_{0}F_{\text{2d}}|$ acting on the hill in the positive $x$-direction, along with an equal and opposite force acting on the flow in the negative $x$-direction. In three dimensions, with a horizontal domain $L_{x}/2,L_{x}/2]\times[-L_{y}/2,L_{y}/2]$, we then define the stress due to a topography $h(x,y)$ as the vector
@@
Note that $\bm{F}_{\text{3d}}$ has units $\mathrm{N}\,\mathrm{m}^{3}\,\mathrm{k}\mathrm{g}^{-1}$ and $\rho_{0}\bm{F}_{\text{3d}}$ has units of force.
In what follows, our domain is as shown in Figure 1, with dimensions $L_{x}\times L_{y}\times H$ and a spatially mean zonal flow $U(t)$ over topography $h(x,y)$. For two-dimensional analysis, we restrict ourselves to the xxzz plane. Moreover, for the ensuing theoretical analysis in this section, we assume $L_{x}$ and $L_{y}$ are large. That is, $L_{x},L_{y}\to\infty$.
@@
Throughout this paper, we use the closely related terms “stress” and “internal wave drag”. To avoid any confusion, we briefly clarify what is meant by each of these terms. By “stress”, or more correctly “form stress”, we mean the total force (scaled by $1/\rho_{0}$) acting on the flow as per (1) or (2). “Internal wave drag” refers to the component of form stress due to the formation of internal waves. Many existing parametrisations focus only on internal wave drag, which is the dominant component of stress in many idealised scenarios, but not if the flow is blocked or sufficiently non-linear.
@@ 2.1 Case 1: Tidal Flow
We consider a periodic tidal flow $U(t)=U_{\text{tidal}}\cos(\omega t)$ in the $x$-direction where $U_{\text{tidal}}$ is a constant and $\omega$ is the tidal frequency. Tidal flows are a major source of internal waves and the stress associated with their generation impacts the strength of both barotropic and internal tides (Ansong et al., 2015; Buijsman et al., 2015). As a result, this flow regime is the main focus in the oceanic literature.
In the work of Bell (1975), linear theory was used to develop an expression for the energy flux generated by internal tidal waves. Bell found his theory to compare favorably with observations, and his parametrisations have since been refined by others including Llewellyn Smith and Young (2002), Khatiwala (2003) and Nycander (2005). Although Bell’s theory was derived for small-scale topography, recent work (Geoffroy et al., 2024; Pollmann and Nycander, 2023) suggests that Bell’s linear theory can still be reliable for moderately large topography, particularly for low mode waves. However, because Bell’s parametrisation and its derivatives provide expressions for the time-averaged energy flux, they cannot be directly converted to time-dependent stress parametrisations.
For our analysis, we focus on two easily applied stress parametrisations closely related to the work of Bell (1975). First, we will discuss an older and ubiquitously applied parametrisation due to Jayne and St. Laurent (2001, hereafter, JSL2001), which is essentially a scaling for stress based on Bell’s formula for the internal tide energy flux. Then, we will analyse a more mathematically rigorous parametrisation due to Shakespeare et al. (2020, hereafter SAH2020). A comparison of other similar parametrisations, with a focus on energy conversion, is given by Green and Nycander (2013).
@@ 2.1.1 Jayne and St Laurent parametrisation
Jayne and St Laurent give a simple estimate for internal wave drag, by assuming that the stress scales in an identical way to the Bell’s energy flux formula. Despite the simple form of the JSL2001 parametrisation, it has been shown to significantly improve the accuracy of tidal models when compared to observations, and is widely used today (Arbic et al., 2018; Ansong et al., 2015). For each grid cell, they set the stress (in each horizontal dimension) to be proportional to $Nh^{2}u$. Here, $N$ is the buoyancy frequency, and $h$ and $u$ are taken to be suitable values of the topography height and tidal velocity in the grid cell. For small $h\ll H$, we can approximate $u$ as the barotropic flow $U(t)$ and $N$ as a constant in the vicinity of the topography. Thus, the total stress over a small hill with root mean square height $h_{rms}$ is estimated by
@@
Here, $\kappa$ is the tunable scaling constant in JSL2001, $N_{B}$ denotes the buoyancy frequency at the seabed, and the subscript “JSL” indicates that this is the drag force due to the JSL2001 theory. In line with our units and scaling for stress in (1) and (2), here $\kappa$ has units of length in three dimensions and is unitless in two dimensions. On the other hand, in JSL2001 the stress is implicitly divided by the horizontal length (in two-dimensions) or area (in three-dimensions) so that $\kappa$ has units of inverse length and may therefore be interpreted as the characteristic wavenumber of the topography. Note that the factor of $1/2$ in (3) could be absorbed into the definition of $\kappa$, but we have included it to more closely match the expression given by JSL2001.
Although simple and widely used, there are some notable limitations of the JSL2001 model. Firstly, the model was only designed for small-scale topography and large ocean depths. In addition:
  1. 1.
    Representing a complicated region of topography by a single wavenumber $\kappa$ is very simplistic and a more sophisticated formulation would be preferred.
  2. 2.
    Using the bottom value of stratification $N_{B}$ has since been shown to be unreliable, and it is more accurate to use an averaged form of the stratification (Zarroug et al., 2010; Lüschow and Von Storch, 2024).
  3. 3.
    The parametrisation (3) is latitude-independent, an omission which Jayne and St Laurent state could be significant when the tidal frequency $\omega$ is close to the Coriolis frequency $f$.
@@ 2.1.2 Shakespeare, Arbic and Hogg parametrisation
In SAH2020, the authors take a more analytic approach to parametrising tidal flow stress, giving an expression which refines that of JSL2001 (Equation (3)). The main novelty of their work is their treatment of “spring” forces, such as bottom-trapped tides, which are out-of-phase with the tidal flow and so do not do any work in the time mean. This is important, because JSL2001 and other prior parametrisations compute stress via a comparison to the time-averaged energy flux, and thus do not account for such spring forces. Since the parametrisation of SAH2020 refines that of JSL2001, including by adding frequency dependence (point 3 above) and a more sophisticated representation of the topographic spectrum (1 above), we mainly compare our numerical simulations with the SAH2020 theory, whilst commenting on the limitations of the JSL2001 theory.
Under standard assumptions, including small topographic height and large ocean depth, SAH2020 give the stress for subcritical latitudes ($|f|<\omega$) and isotropic topography as
@@
Here, $\alpha\in\{0,1\}$ is such that
Then, $\hat{h}$ denotes the Fourier transform of $h$ in $t$, $x$ and $y$ to transformed variables $\omega$, $k$ and $\ell$ respectively with $K\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sqrt{k^{2}+\ell^{2}}$. Note that we use the Fourier transform convention
@@
Since we are assuming the topography is isotropic (and independent of time), $\hat{h}$ is a function of the total wavenumber $K$ only, as indicated in (4). We also remark that (4) is only the $x$-component of the stress, as this is the primary direction of our flow and focus of our analysis.
As a direct comparison to the JSL2001 parametrisation, a clear frequency dependency has been added and the constant $\kappa$ in (3) has been replaced with the analytic expression
$\frac{1}{2\pi h_{\text{rms}}^{2}}\int_{0}^{\infty}|\hat{h}(K)|^{2}K^{2}\mathrm{d}K.$
In particular, in the limit where $N\gg\omega\gg f$, $F_{\text{3d}}$ reduces exactly to the JSL2001 expression (3), albeit with an analytic expression for what $\kappa$ should be set to.
In SAH2020, a more general expression than (4) is stated which accounts for wave reflections off the ocean surface. However, such reflections are difficult to parametrise (e.g., Shakespeare et al., 2021b), and for a narrow, isolated hill in a deep ocean the number of waves reflecting back onto the topography is small. Therefore, we omit any reflection terms from (4) for simplicity. Nevertheless, we note that these reflections become important even for an isolated hill when the half-vertical wavelength $\lambda_{v}$ compared to the ocean depth $H$, known as the mode-number
$n=\frac{H}{\lambda_{v}/2}=\frac{KH\sqrt{N^{2}-\alpha\omega^{2}}}{\pi\sqrt{\omega^{2}-f^{2}}},$ (7)
approaches 1 (from above). As a result, we expect (4) may have deviations from the actual stress for small $K$, $H$ or $N$.
In the case of an isolated Gaussian hill with width $W$, whereby
@@
Importantly, (9) suggests that the stress depends linearly on the hill width $W$, at least for Gaussian hills.
The expression for supercritical (bottom-trapped) latitudes $|f|\geq\omega$ is similar, but with the stress out of phase with velocity. In particular, for a Gaussian hill
@@
$\ \times\begin{cases}\sqrt{(N^{2}-\alpha\omega^{2})(\omega^{2}-f^{2})}\,\cos(\omega t),\ |f|<\omega,\\ \sqrt{(N^{2}-\alpha\omega^{2})(f^{2}-\omega^{2})}\,\sin(\omega t),\ |f|\geq\omega,\end{cases}$ (10)
where the subscript SAH3d indicates that this is the three-dimensional stress due to the SAH2020 theory. In a subsequent paper (Shakespeare et al., 2021a), the authors also give a parametrisation for two dimensions. Namely, for a Gaussian hill $h(x)=h_{0}\,e^{-x^{2}/(2W^{2})}$,
@@
$\ \times\begin{cases}\sqrt{(N^{2}-\alpha\omega^{2})(\omega^{2}-f^{2})}\cos(\omega t),&|f|<\omega,\\ \sqrt{(N^{2}-\alpha\omega^{2})(f^{2}-\omega^{2})}\sin(\omega t),&|f|\geq\omega.\end{cases}$ (11)
Note that (2.1.2) is identical to (2.1.2) except without the factor of $\pi\sqrt{\pi}W/4$. To obtain (2.1.2) and (2.1.2) one also needs to assume that $N$ is constant, which we will do throughout our analysis. As discussed before, in the more physical setting where $N$ is varying function of depth, it is better to use an averaged value of $N$, at least for sufficiently large hill widths (Lüschow and Von Storch, 2024).